OpenStudy (anonymous):

Help me with a 45-45-90 triangle? I need someone to walk me through/teach me how to solve this problem..

OpenStudy (anonymous):

Solve for x and y. It's a 45-45-90 triangle, and one leg is sqr rt 2.I think x (the other leg ) would then be square root 2 but I am not sure

OpenStudy (anonymous):

in 45-45-90 triangle 2 sides are same and third one is different. Sides opposite 45-45 are same.

OpenStudy (anonymous):

It looks like this @zepdrix

OpenStudy (anonymous):

if sides opposite 45-45 are X , then hypotenuse is X sq.rt 2

zepdrix (zepdrix):



OpenStudy (anonymous):

@zepdrix yup that's what it looks like @rukhsar What do you mean X square root 2?

OpenStudy (anonymous):

\[\sqrt{2}\] X

OpenStudy (anonymous):

Does that mean x (the other leg) would be \[\sqrt{2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Y = sq.rt 2 multiply by sq.rt 2

OpenStudy (anonymous):

it's the property of 45-45-90 triangle

OpenStudy (anonymous):

\[\sqrt{2} \sqrt{2}\]

OpenStudy (anonymous):

so Y= 2

OpenStudy (anonymous):

so \[\sqrt{2} multiply \sqrt{2} would that equal \sqrt{4}\]

OpenStudy (anonymous):

yes and sq.rt 4 = 2

OpenStudy (anonymous):

Ah! I see, thank you so much! I was really struggling understanding it, I think I understand it now!

OpenStudy (anonymous):

np